OpenStudy (anonymous):

Multiply and simplify this expression: (√2 + √5)(√2 - √5)

OpenStudy (anonymous):

2 - 5

OpenStudy (anonymous):

a^2 - b^2 = (a+b) (a - b)

OpenStudy (anonymous):

2 - 5 = -3